OpenStudy (anonymous):

Whats the value of this?? Trigonometric options A)pi/4 B)-pi/4 C)pi/2 D)-pi/2 tan^−1 (x/y)−tan^−1 (x−y/x+y) =

OpenStudy (anonymous):

\[\tan^{-1} (x/y)- \tan^{-1} (x-y/x+y) \]

OpenStudy (anonymous):

x,y>0

OpenStudy (anonymous):

@abb0t @hartnn @mathaddict4471

OpenStudy (anonymous):

@Loser66 @CGGURUMANJUNATH

OpenStudy (abb0t):

y ≠ 0, x = 0

OpenStudy (anonymous):

options are \[\pi/4..............(-\pi/4) ..........\pi/2 .............(-\pi/2)\]

OpenStudy (anonymous):

@agent0smith @BulletWithButterflyWings @Hero @ladydeadpool @mathaddict4471 @phi @primeralph @Psymon @Preetha

OpenStudy (anonymous):

@amistre64

OpenStudy (primeralph):

You don't have an equality sign anywhere.

OpenStudy (tkhunny):

I presume you mean (x-y)/(x+y) and not what you have written?

OpenStudy (anonymous):

the above equals which option???

OpenStudy (primeralph):

@cambrige You were asked a question by @tkhunny

OpenStudy (anonymous):

No

OpenStudy (anonymous):

the answer to @tkhunny is no :)

OpenStudy (tkhunny):

You mean \(\left(x - \dfrac{y}{x} + y\right)\), exactly as you have it written?

OpenStudy (tkhunny):

Well, \(atan\left(\dfrac{x}{y}\right)-atan\left(\dfrac{x-y}{x+y}\right) = \dfrac{\pi}{4}\), but you go ahead and work the other problem that has no such value.

OpenStudy (tkhunny):

It may be simpler to show this: \(\sin\left(atan\left(\dfrac{x}{y}\right)-atan\left(\dfrac{x-y}{x+y}\right)\right) = \dfrac{1}{\sqrt{2}}\) The result requested follows immediately.

OpenStudy (anonymous):

whats atan??

OpenStudy (anonymous):

it is how u have put it @tkhunny

OpenStudy (tkhunny):

How about that. It happens. :-) \(atan(x)\) is, for most purposes, exactly the same as \(tan^{-1}(x)\). It's just easier to write. One day, you may be concerned about a fundamental difference between the two, but I'm guessing that day is not today.